Outback Outfitters sells recreational equipment. One of the companys products, a small camp stove, sells for $110 per unit.
Answer #1

Answer to Requirement 1.

Unit Selling Price = $110
Unit Variable Expense = $77
Unit Contribution Margin = Unit Selling Price - Unit Variable Expense
Unit Contribution Margin = $110 - $77
Unit Contribution Margin = $33

Break Even Sales (in Units) = Fixed Expense / Unit Contribution Margin
Break Even Sales (in Units) = $161,700 / $33
Break Even Sales (in Units) = 4,900 Units

Break Even Sales (in Dollar) = Break Even Sales (in Units) * Unit Selling Price
Break Even Sales (in Dollar) = 4,900 * $110
Break Even Sales (in Dollar) = $539,000

Answer to Requirement 2.

Higher Break Even Point

Increase in Variable Expenses per stove as percentage of selling price will decrease Contribution Margin per Stove. Decrease in Contribution Margin per stove will result in lower Contribution Margin ratio. The decrease in Contribution Margin will result in Higher Break Even point.

Answer to Requirement 3.

Current Unit Selling price = $110
Proposed Unit Selling Price = $110 – ($110 * 10%) = $99

Current Units sold = 19,000 Stoves
Proposed Unit Sold = 19,000 + (19,000 * 25%) = 23,750 stoves

Sales Variable Expenses Contribution Margin Fixed Expenses Net Operating Income Outback Outfitters Present 19,000 Stoves Tota

Answer to Requirement 4.

Proposed Unit Contribution Margin = $99 - $77
Proposed Unit Contribution Margin = $22

Desired Unit Sales = (Fixed Cost + Target Profit) / Unit Contribution Margin
Desired Unit Sales = ($161,700 + $72,000) / $22
Desired Unit Sales = 10,623 Stoves